A student, sitting on a stool rotating at a rate of 26 rev/min, holds masses in each hand. When his arms are extended, the total rotational inertia of the system is 8.8 kg · m2. He pulls his arms in close to his body, reducing the total rotational inertia to 5.1 kg ·m2 . External torques are negligible. The new rotational speed of the system is
Answer #1

Answer-

Conservation of angular momentum

I*ω = constant.

I1*ω1 =I2*ω2

8.8 * 26 = 5.1 * ω2

ω2 = 44.86 rev/min.

new rotational speed of the system is ω2 = 44.86 rev/min.